# Amortized SHAP values via sparse Fourier function approximation

### v1

@@ B.2 Sparse and low-degree Fourier transform algorithms
Under the additional assumption that $h$ is $k$-sparse, works such as Cheraghchi & Indyk (2017); Amrollahi et al. (2019); Scheibler et al. (2015); Kushilevitz & Mansour (1993); Li & Ramchandran (2015); Li et al. (2015) provide algorithms that obtain the Fourier transform more efficiently. In particular, Amrollahi et al. (2019) provide algorithms with query complexity $O(nk)$ and time complexity $O(nk\log k)$ time. Assuming further that the function is of degree $d=o(n)$, the query complexity reduces to $O(kd\log n)$, with run time still polynomial in $n,k,d$. Crucially, even if the function $h$ is not $k$-sparse, Algorithm RobustSWHT of Amrollahi et al. (2019) yields the best $k$-sparse approximation in the $\ell_{2}-\ell_{2}$ sense. More precisely, let us denote by $h_{k}:\{0,1\}^{n}\rightarrow$ the function that is formed by only keeping the top $k$ non-zero Fourier coefficients of $h$ and setting the rest to zero. Then the algorithm returns a $O(k)$-sparse function $g$ such that:
@@
B.3 Shapley values in the context of Machine learning
@@
In the context of ML, many works have derived a different notion of Shapley value depending on what they mean by data distribution, deleted features, etc. We refer the reader to the survey by Sundararajan & Najmi (2020); Janzing et al. (2020) for a comprehensive overview. In this work we focus on the notion of SHAP introduced by Lundberg & Lee (2017); Lundberg et al. (2020) also known as “Interventional” (Janzing et al., 2020; Van den Broeck et al., 2021) or “Baseline” SHAP (Sundararajan & Najmi, 2020) where the missing features are integrated out from the marginal distribution as opposed to the conditional distribution see Section 2. As pointed out by Janzing et al. (2020) there are two main ways to define the SHAP value “interventional” and “observational” SHAP. These are referred to by Sundararajan & Najmi (2020) as “baseline” and “conditional” SHAP respectively.
@@
As pointed out by Janzing et al. (2020) the difference between these definitions can be better viewed with the lens of causality (Pearl, 2009). Roughly speaking “observational” SHAP tells us about how influential a feature is to the prediction of the predictor if it goes from the state of being unobserved to observed. “Interventional” SHAP is causal and tells us how influential a feature is if we were to reach in (through a process called an intervention) and change that feature in order to change the prediction. Put into the context of credit scores and loan approvals, “observational” SHAP will provide us with important features which are “observed” by the predictor and hence are influential in predicting if a particular loan request will be rejected or approved. Interventional SHAP would tell us which feature we could change or “intervene” in order to change the outcome of the loan request.
@@
The original (ML) SHAP paper (Lundberg & Lee, 2017) proposes “observational” SHAP as the correct notion of SHAP. Van den Broeck et al. (2021); Arenas et al. (2021) provide intractability results for observational SHAP in a variety of simple distributional assumptions on the data and simple predictors $f$. This has lead to many attempts to approximate observational SHAP values (Lundberg et al., 2020; Covert & Lee, 2020; Aas et al., 2021; Kwon et al., 2021; Sundararajan & Najmi, 2020). It is interesting to note that the version of “Kernel”-SHAP in Lundberg & Lee (2017) is also an approximation for observational SHAP values that ends up coinciding precisely with interventional SHAP values, which explains a lot of the confusion in the community. Janzing et al. (2020) boldly claims that researchers should stop the pursuit of approximations to “observation” SHAP values as it lacks certain properties, for example, sensitivity i.e. the SHAP value of a feature can be non-zero while the predictor $f$ has no dependence on that feature. This phenomenon happens because when features are correlated, the presence of a feature can provide information about other features that the predictor does depend on. This does not happen in interventional SHAP. Finally, Chen et al. (2020) argues that both SHAP definitions are worthy of pursuit. They emphasize that the interventional framework provides explanations that are more “true to the data”, and the observational approach’s explanations are more “true to the model”.
Appendix C Proofs
@@ C.1 Proof of propositions
Where Equation C.1 holds because $g$ is only dependent on the variables in $S$ and Equation C.1 holds by checking the inner sum has an equal number of $1$ and $-1$ added together.
@@ C.2 Proof of Lemma 1
One can see by checking for different values of $x_{i},x_{i}^{*},f_{i}\in\{0,1\}$ that $(-1)^{f_{i}x^{*}_{i}}-(-1)^{f_{i}x_{i}}=2f_{i}(x_{i}-x_{i}^{*})$. To determine $(-1)^{\langle f_{-i},x^{*}_{S}\oplus x_{[n]\setminus S\cup\{i\}}\rangle}$, we partition $[n]\setminus\{i\}$ into two subsets $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$ and $B\triangleq[n]\setminus A\cup\{i\}$. Doing this, we can factor out $(-1)^{\langle f_{-i},x_{-i}\rangle}$ and determine the rest of the sign based on the number of indices in $S$ where $x$ and $x^{*}$ disagree and $f_{i}=1$. This is equal to $|A\cap S|$:
@@
In Equation C.2, we use the following identity $\binom{n-1}{a+b}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}=\binom{n-1}{|A|}\binom{|A|}{a}\binom{n-|A|-1}{b}$ which can be checked algebraically by the reader by simply writing down each binomial term as factorials and doing the cancellations.
In Equation C.2, we use the following identity $\sum_{b=0}^{n-|A|-1}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}=\binom{n}{|A|+1}$ which holds because of the following double-counting argument. We want to pick $|A|+1$ items out of $n$ items numbered $1,\dots,n$. Say we pick $|A|+1$ items and order them according to their number. We condition on the $(a+1)^{th}$ chosen item (after ordering). The $(a+1)^{th}$ chosen item’s number can be equal to $(a+b+1)$, where $b$ ranges from 00 to $n-|A|-1$. Choosing the preceding $a$ and trailing $|A|-a$ items can be done in $\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$ ways.
Based on Equation 8, we see that the innermost summation in Equation C.2 is only dependent on $|A|$. Thus, we rewrite Equation C.2 as follows:
@@ Appendix D Datasets
Poelwijk et al. (2019) study the fluorescence brightness of all $2^{13}$ distinct variants of the Entacmaea quadricolor fluorescent protein, mutated at $13$ different sites.
@@
Nugteren & Codreanu (2015) measures the running time of a matrix product using a parameterizable SGEMM GPU kernel, configured with different parameter combinations. The input has $14$ categorical features. After one-hot encoding the dataset is $40$-dimensional.
@@
Wu et al. (2016) study the “fitness” of variants of protein GB1, that are mutated at four different sites. Fitness, in this work, is a quantitative measure of the stability and functionality of a protein variant. Given the $20$ possible amino acids at each site, they report the fitness for $00020^{4}=160,$ 000 possible variants, which we represent with one-hot encoded 80-dimensional binary vectors.
@@
Sarkisyan et al. (2016) estimate the fluorescence brightness of random mutations over the green fluorescent protein sequence of Aequorea victoria (avGFP) at 236 amino acid sites. We transform the amino acid features into binary features indicating the absence or presence of a mutation at each amino acid site. This converts the original $54,024$ distinct amino acid sequences of length 236 into $49,089$ 236-dimensional binary data points.
Appendix E FourierShap Implementation
We implemented four versions of our algorithm, i.e., Equation 5, using Google JAX library (Bradbury et al., 2018). JAX provides a flexible framework for developing high-performace functions for vectorized computations. JAX enables automatic performance optimisation of algebraic computation as well as just-in-time (JIT) compilation of the vectorized functions for faster iterations at runtime.
@@ Appendix E FourierShap Implementation
Appendix F Experiment Details
The code for running the experiments and the implementations of all modules will be open-sourced once the double-blind review process is over. We run all experiments on a machine with one NVIDIA GeForce RTX 4090 GPU.
F.1 Black-box
For the Entacmaea and SGEMM datasets, we train fully connected neural networks with 3 hidden layers containing 300 neurons each. The network is trained using the means-squared loss and ADAM optimizer with a learning rate of 0.01. For GB1 we train ensembles of trees models of varying depths using the random forest algorithm using the sklearn library (Pedregosa et al., 2011). For avGFP we train again, ensembles of trees models with 10 trees of varying depths using the cat-boost algorithm/library(Dorogush et al., 2018). All other setting are set to the default in both cases. Model accuracy for different depths are plotted in Figure 4.
@@ F.1 Black-box
F.1.1 FourierSHAP
For the first step of the FourierSHAP, i.e., computing a sparse Fourier approximation of the black-box model, we use an implementation of a sparse Walsh-Hadamard Transform (sparse WHT) a.k.a Fourier transform algorithm (Amrollahi et al., 2019) for each of the four trained models. The implementation is done using Google JAX (Bradbury et al., 2018) library as part of this work, and is available in the project repository.
For the second step of FourierSHAP, i.e., utilizing the approximated Fourier spectrum to compute SHAP values as per Equation 5, we use our “Precompute” implementation. We discussed the details in the previous Section, Section E.
@@ F.1.1 FourierSHAP
F.1.2 KernelShap
For KernelShap, we use the standard library provided by the writers of the paper (Lundberg & Lee, 2017)  [^3] https://github.com/slundberg/shap with its default settings. As part of the default setting, the paired sampling trick is also enabled, which shown to be beneficial for faster convergence of KernelSHAP. KernelShap is written in C and is to our knowledge the fastest implementation of this algorithm.
@@ F.1.3 Fast-SHAP
  • Imputer: ”Imputer” is a FastSHAP module used in the computation of neural networks’ loss, acting as the value function in SHAP formula, that generates model’s prediction using a strategy for treating features excluded in the subset, given the predictor and a subset of features. To compute interventional SHAP values, we use MarginalImputer, implemented in the original FastSHAP repo, which computes mean predictions when using the backgorund dataset’s values for excluded features. Therefore, each trained FastSHAP model is specific to a fixed background dataset, as the Imputer used in its loss is. In our experiments, we use four different background datasets with multiple sizes per (real-world) dataset, which lead to training four FastSHAP models per setting.
@@
F.1.4 LinRegShap
LinRegShap, is a variance-reduced version of KernelSHAP (Covert & Lee, 2020). We again use the implementation of the original writers [^4] https://github.com/iancovert/shapley-regression. Their implementation includes automatic detection of the convergence of stochastic sampling which is meant to speed up the algorithm by taking less samples from the black box. Furthermore, as per the default setting, we also allowed paired sampling.
@@ F.1.4 LinRegShap
F.1.5 DeepLift
For DeepLift we as well use the library of Lundberg & Lee (2017)  [^5] https://github.com/slundberg/shap with default settings. In this setting the neural network is passed to the algorithm on a GPU to make sure this algorithm is as fast as possible.
F.2 Trees

### v2

@@ A.2 Sparse and low-degree Fourier transform algorithms
Under the additional assumption that $h$ is $k$-sparse, works such as Cheraghchi and Indyk [2017], Amrollahi et al. [2019], Scheibler et al. [2015], Kushilevitz and Mansour [1993], Li and Ramchandran [2015], Li et al. [2015] provide algorithms that obtain the Fourier transform more efficiently. In particular, Amrollahi et al. [2019] provide algorithms with query complexity $O(nk)$ and time complexity $O(nk\log k)$ time. Assuming further that the function is of degree $d=o(n)$, the query complexity reduces to $O(kd\log n)$, with run time still polynomial in $n,k,d$. Crucially, even if the function $h$ is not $k$-sparse, Algorithm RobustSWHT of Amrollahi et al. [2019] yields the best $k$-sparse approximation in the $\ell_{2}-\ell_{2}$ sense. More precisely, let us denote by $h_{k}:\{0,1\}^{n}\rightarrow$ the function that is formed by only keeping the top $k$ non-zero Fourier coefficients of $h$ and setting the rest to zero. Then the algorithm returns a $O(k)$-sparse function $g$ such that:
@@
A.3 Shapley values in the context of Machine learning
@@
In the context of ML, many works have derived a different notion of Shapley value depending on what they mean by data distribution, deleted features, etc. We refer the reader to the survey by Sundararajan and Najmi [2020], Janzing et al. [2020] for a comprehensive overview. In this work we focus on the notion of SHAP introduced by Lundberg and Lee [2017], Lundberg et al. [2020] also known as “Interventional” [Janzing et al., 2020, Van den Broeck et al., 2021] or “Baseline” SHAP [Sundararajan and Najmi, 2020] where the missing features are integrated out from the marginal distribution as opposed to the conditional distribution (see Section 2). As pointed out by Janzing et al. [2020] there are two main ways to define the SHAP value “interventional” and “observational” SHAP. These are referred to by Sundararajan and Najmi [2020] as “baseline” and “conditional” SHAP respectively.
@@
As pointed out by Janzing et al. [2020] the difference between these definitions can be better viewed with the lens of causality [Pearl, 2009]. Roughly speaking “observational” SHAP tells us about how influential a feature is to the prediction of the predictor if it goes from the state of being unobserved to observed. “Interventional” SHAP is causal and tells us how influential a feature is if we were to reach in (through a process called an intervention) and change that feature in order to change the prediction. Put into the context of credit scores and loan approvals, “observational” SHAP will provide us with important features which are “observed” by the predictor and hence are influential in predicting if a particular loan request will be rejected or approved. Interventional SHAP would tell us which feature we could change or “intervene” in order to change the outcome of the loan request.
@@
The original (ML) SHAP paper [Lundberg and Lee, 2017] proposes “observational” SHAP as the correct notion of SHAP. Van den Broeck et al. [2021], Arenas et al. [2021] provide intractability results for observational SHAP in a variety of simple distributional assumptions on the data and simple predictors $f$. This has lead to many attempts to approximate observational SHAP values [Lundberg et al., 2020, Covert and Lee, 2020, Aas et al., 2021, Kwon et al., 2021, Sundararajan and Najmi, 2020]. It is interesting to note that the version of “Kernel”-SHAP in Lundberg and Lee [2017] is also an approximation for observational SHAP values that ends up coinciding precisely with interventional SHAP values, which explains a lot of the confusion in the community. Janzing et al. [2020] boldly claims that researchers should stop the pursuit of approximations to “observation” SHAP values as it lacks certain properties, for example, sensitivity i.e. the SHAP value of a feature can be non-zero while the predictor $f$ has no dependence on that feature. This phenomenon happens because when features are correlated, the presence of a feature can provide information about other features that the predictor does depend on. This does not happen in interventional SHAP. Finally, Chen et al. [2020] argues that both SHAP definitions are worthy of pursuit. They emphasize that the interventional framework provides explanations that are more “true to the model”, and the observational approach’s explanations are more “true to the data”.
Appendix B Proofs
@@ B.1 Proof of propositions
Where Equation B.1 holds because $g$ is only dependent on the variables in $S$ and Equation B.1 holds by checking the inner sum has an equal number of $1$ and $-1$ added together.
@@ B.2 Proof of Lemma 1
By checking all 8 possible combinations of $x_{i},x_{i}^{*},f_{i}\in\{0,1\}$, one can see that $(-1)^{f_{i}x^{*}_{i}}-(-1)^{f_{i}x_{i}}=2f_{i}(x_{i}-x_{i}^{*})$. This is simply because the two exponents only differ when $f_{i}=1$ and $x_{i}\neq x_{i}^{*}$.
To determine $(-1)^{\langle f_{-i},x^{*}_{S}\oplus x_{[n]\setminus S\cup\{i\}}\rangle}$, we partition $[n]\setminus\{i\}$ into two subsets $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$ and $B\triangleq[n]\setminus A\cup\{i\}$. Doing this, we can factor out $(-1)^{\langle f_{-i},x_{-i}\rangle}$ and determine the rest of the sign based on the number of indices in $S$ where $x$ and $x^{*}$ disagree and $f_{i}=1$. This is equal to $|A\cap S|$:
@@
In Equation B.2, we use the following identity: $\binom{n-1}{a+b}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}=\binom{n-1}{|A|}\binom{|A|}{a}\binom{n-|A|-1}{b}$. This can be checked algebraically by simply writing down each binomial term as factorials and doing the cancellations:
$\binom{n-1}{a+b}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$
$=\frac{(n-1)!}{(a+b)!(n-a-b-1)!}\cdot\frac{(a+b)!}{a!b!}\cdot\frac{(n-a-b-1)!}{(|A|-a)!(n-a-b-1-(|A|-a))!}$
$=\frac{(n-1)!}{a!b!(|A|-a)!(n-|A|-b-1)!}.$
$=\frac{(n-1)!}{|A|!(n-|A|-1)!}\cdot\frac{|A|!}{a!(|A|-a)!}\cdot\frac{(n-|A|-1)!}{b!(n-|A|-1-b)!}$
$=\binom{n-1}{|A|}\binom{|A|}{a}\binom{n-|A|-1}{b}$
In Equation B.2, we use $\sum_{b=0}^{n-|A|-1}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}=\binom{n}{|A|+1}$ which holds because of the following double-counting argument. The term $\binom{n}{|A|+1}$ counts the number of ways to choose a subset of size $|A|+1$ from a set of $n$ elements. Imagine elements are numbered from 1 to n, and $\pi_{1}<\pi_{2}<...<\pi_{|A|+1}\in[n]$ represent $|A|+1$ chosen elements. Let’s condition on $\pi_{a+1}=(a+b+1)$ where possible values for $b$ can only be $10\leq b\leq n-|A|-$ 1. This is due to the fact that $\pi_{a+1}<a+1$ implies $\pi_{1},...,\pi_{a}$ are chosen from less than $a$ elements, and similarly $\pi_{a+1}>n-|A|+a$ implies $\pi_{a+2},...,\pi_{|A|+1}$ ($|A|-a$ chosen elements) are chosen from less than $|A|-a$ elements, which are both impossible. Given the condition, $a$ elements numbered lower than $(a+b+1)$ and $|A|-a$ elements numbered larger than $(a+b+1)$ are also chosen. This is possible in $\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$ ways. Therefore, keeping in mind that $a$ is fixed, $\sum_{b=0}^{n-|A|-1}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$ should give the number of ways to choose a subset of size $|A|+1$ from a set of $n$ elements, through the new perspective.
Based on Equation 8, we see that the innermost summation in Equation B.2 is only dependent on $|A|$. Thus, we rewrite Equation B.2 as follows:
@@ Appendix C Datasets
Poelwijk et al. [2019] study the fluorescence brightness of all $2^{13}$ distinct variants of the Entacmaea quadricolor fluorescent protein, mutated at $13$ different sites.
@@
Nugteren and Codreanu [2015] measures the running time of a matrix product using a parameterizable SGEMM GPU kernel, configured with different parameter combinations. The input has $14$ categorical features. After one-hot encoding the dataset is $40$-dimensional.
@@
Wu et al. [2016] study the “fitness” of variants of protein GB1, that are mutated at four different sites. Fitness, in this work, is a quantitative measure of the stability and functionality of a protein variant. Given the $20$ possible amino acids at each site, they report the fitness for $00020^{4}=160,$ 000 possible variants, which we represent with one-hot encoded 80-dimensional binary vectors.
@@
Sarkisyan et al. [2016] estimate the fluorescence brightness of random mutations over the green fluorescent protein sequence of Aequorea victoria (avGFP) at 236 amino acid sites. We transform the amino acid features into binary features indicating the absence or presence of a mutation at each amino acid site. This converts the original $54,024$ distinct amino acid sequences of length 236 into $49,089$ 236-dimensional binary data points.
Appendix D FourierShap Implementation
We implemented four versions of our algorithm, i.e., Equation 5, using Google JAX library [Bradbury et al., 2018]. JAX provides a flexible framework for developing high-performace functions for vectorized computations. JAX enables automatic performance optimisation of algebraic computation as well as just-in-time (JIT) compilation of the vectorized functions for faster iterations at runtime.
@@ Appendix D FourierShap Implementation
Appendix E Experiment Details
The code for running the experiments and the implementations of all modules will be open-sourced once the double-blind review process is over. We run all experiments on a machine with one NVIDIA GeForce RTX 4090 GPU, on servers with Intel(R) Xeon(R) CPU E3-1284L v4 @ 2.90GHz, restricting the memory/RAM to 20 GB, which was managed with Slurm.
E.1 Black-box
For the Entacmaea and SGEMM datasets, we train fully connected neural networks with 3 hidden layers containing 300 neurons each. The network is trained using the means-squared loss and ADAM optimizer with a learning rate of 0.01. For GB1 we train ensembles of trees models of varying depths using the random forest algorithm using the sklearn library [Pedregosa et al., 2011]. For avGFP we train again, ensembles of trees models with 10 trees of varying depths using the cat-boost algorithm/library[Dorogush et al., 2018]. All other setting are set to the default in both cases. Model accuracy for different depths are plotted in Figure 4.
@@ E.1 Black-box
E.1.1 FourierSHAP
For the first step of the FourierSHAP, i.e., computing a sparse Fourier approximation of the black-box model, we use an implementation of a sparse Walsh-Hadamard Transform (sparse WHT) a.k.a Fourier transform algorithm [Amrollahi et al., 2019] for each of the four trained models. The implementation is done using Google JAX [Bradbury et al., 2018] library as part of this work, and is available in the project repository.
For the second step of FourierSHAP, i.e., utilizing the approximated Fourier spectrum to compute SHAP values as per Equation 5, we use our “Precompute” implementation. We discussed the details in the previous Section, Section D.
@@ E.1.1 FourierSHAP
E.1.2 KernelShap
For KernelShap, we use the standard library provided by the writers of the paper [Lundberg and Lee, 2017]  [^3] https://github.com/slundberg/shap with its default settings. As part of the default setting, the paired sampling trick is also enabled, which shown to be beneficial for faster convergence of KernelSHAP. KernelShap is written in C and is to our knowledge the fastest implementation of this algorithm.
@@ E.1.3 Fast-SHAP
  • Imputer: "Imputer" is a FastSHAP module used in the computation of neural networks’ loss, acting as the value function in SHAP formula, that generates model’s prediction using a strategy for treating features excluded in the subset, given the predictor and a subset of features. To compute interventional SHAP values, we use MarginalImputer, implemented in the original FastSHAP repo, which computes mean predictions when using the backgorund dataset’s values for excluded features. Therefore, each trained FastSHAP model is specific to a fixed background dataset, as the Imputer used in its loss is. In our experiments, we use four different background datasets with multiple sizes per (real-world) dataset, which lead to training four FastSHAP models per setting.
@@
Table 1 shows the time required to train the FastSHAP models used in our black-box experiments. Both FourierShap (ours) and FastSHAP are amortized methods that have a heavier “pre-computation” step. For FastShap this pre-computation appears as training an MLP that directly predicts SHAP values and for us this appears as computing a Fourier transform. When comparing FastShap’s initial training time to ours (reported in Figure 1), we can see our method has a considerably lower pre-computation time.
Dataset Black-box model FastSHAP size Training time (m) for background dataset size Total training time (m)
10 20 30 40
Entacmaea MLP Small 5 8 15 18 46
Medium 2 4 6 10 22
Large 3 6 11 11 31
SGEMM MLP Small 47 130 249 173 599
Medium 55 52 162 128 397
Large 72 73 152 91 388
GB1 Random Forest (depth=5) Small 114 148 42 173 477
Medium 83 207 55 130 475
Large 54 102 155 184 495
avGFP Catboost (depth=5) Small 24 28 36 39 127
Medium 35 58 41 62 196
Large 11 20 22 25 78
Table 1: Training times of FastSHAP models used in our black-box experiments (in minutes). Unlike FourierSHAP (ours), FastSHAP needs to be separately trained for each background dataset. Part of the difference in the training times are due to the variance in the number of training epochs before the early stopping occurs. Original datasets are also of varying sizes that lead to different number of training samples per epoch.
E.1.4 LinRegShap
LinRegShap, is a variance-reduced version of KernelSHAP [Covert and Lee, 2020]. We again use the implementation of the original writers [^4] https://github.com/iancovert/shapley-regression. Their implementation includes automatic detection of the convergence of stochastic sampling which is meant to speed up the algorithm by taking less samples from the black box. Furthermore, as per the default setting, we also allowed paired sampling.
@@ E.1.4 LinRegShap
E.1.5 DeepLift
For DeepLift we as well use the library of Lundberg and Lee [2017]  [^5] https://github.com/slundberg/shap with default settings. In this setting the neural network is passed to the algorithm on a GPU to make sure this algorithm is as fast as possible.
E.2 Trees